EDITORIAL

Terry Goodman

In recent years the Department of Mathematics and Computer Science at CMSU has initiated a number of programs/projects designed to aid in recruiting, encouraging, and rewarding outstanding mathematics students.

In recruiting, we believe it is important to identify and contact students early – before they come here to the university. We currently sponsor a high school mathematics competition and co–sponsor a junior high competition each spring. We are attempting to capitalize on these students’ interest in mathematics and we try to make them aware of the programs and opportunities found in our department. Further, through these competitions, we are able to get to know a number of junior high and high school mathematics teachers. We encourage these teachers to provide us with the names of students who demonstrate ability and interest in mathematics.

A number of our faculty participate in activities that bring them in contact with high school mathematics students. Most notably, several faculty participate in the MAA High School Lectureship Program which allows them to connect with students who may be looking toward a career in mathematics or a related area.

We believe it is very important to recognize and reward outstanding mathematics students. We currently sponsor a variety of activities during Mathematics Awareness Week each April. Among these activities is a banquet where outstanding high school mathematics students and their teachers are recognized. Also, at this banquet a number of our current mathematics majors and minors are awarded scholarships. We find that often the high school students recognized at this banquet later become some of our scholarship recipients.

We recognize that our efforts to recruit and reward students are certainly not unique. We would like to hear from our readers concerning their efforts in this area. We will then be able to share this information through future issues of MJMS.